# A Comprehensive Review of Few-shot Action Recognition

### v1

@@ IV-B3 Discussion
IV-C Category Prototype Learning
@@ IV-C1 Aggregation-based
Text-Guided. Furthermore, several methods are inspired by the strong generalization ability of textual information to modulate the action prototypes with the guidance of semantic priors to resolve the challenge of inaccurate prototype estimation caused by data scarcity, as illustrated in Figure 6(d). For instance, Ni et al. [94] extract semantic features from label texts with a frozen CLIP [97] text encoder and further enhanced by a semantic-enhanced module. The prototypes are enhanced by semantic features with flexible techniques such as weighted average and multi-head attention. Similarly, Wang et al. [93] adaptively refine visual prototypes with transferable textual concepts from CLIP, which is achieved by implementing a temporal Transformer to fuse the textual and visual features in the support set. These approaches demonstrate how integrating multimodal semantic information can significantly improve the fidelity and effectiveness of prototype representations.
IV-C2 Learning-based
@@
IV-C3 Compound Methods
Due to the intertwining of temporal and spatial information in videos, alongside the potential occurrence of actions at diverse speeds, a single approach to prototype calculation may overlook essential video details. Consequently, certain methodologies advocate for computing compound prototypes to fortify the resilience of prototype representations. For instance, Huang et al. [139] summarize each video into compound prototypes consisting of a group of global prototypes and a group of focused prototypes, and then compare video similarity based on the compound prototypes. Each global prototype summarizes a specific aspect of the entire video, while focused prototypes target certain timestamps, with bipartite matching applied to handle temporal variations in actions. The authors further introduce a method to learn prototypes combined with three types of prototypes [140]. The first type describes specific aspects of the action regardless of its timestamp. The second type focuses on specific timestamps and uses bipartite matching to handle temporal variations. The third type is generated from the timestamp-centered prototypes, regularizing temporal consistency and providing an auxiliary summarization of the entire video. Jiang et al. [141] introduce a dual-prototype network that integrates class-specific and query-specific attentive learning. The class-specific attentive learning approach enhances prototype representativeness by calculating within-class similarity for each support sample, thereby reducing the influence of noise and outliers. In contrast, the query-specific attention mechanism adaptively reweights support samples based on their similarities to each query sample, which increases inter-class discrimination by enlarging differences among support samples. In summary, these methodologies highlight the effectiveness of compound prototype representations in bolstering prototype robustness significantly and enhancing the ability to generalize to novel categories.
IV-C4 Discussion
TABLE III: Results of different Category Prototype Learning methods.
@@ IV-C4 Discussion
The objective of category prototype learning in few-shot action recognition is to learn representative and generalized action representations for novel categories from a few samples, thereby achieving stable recognition performance in each meta-task. The conventional average-based method may be affected by sample capacity and extreme samples, making mean-of-class prototypes inadequate in representing the distribution of samples. As shown in Table III, the prototype learning methods designed for FSAR have achieved significant performance. More recent approaches that enhance class prototype representations with semantic prior knowledge of large models have demonstrated significant performance benefits across multiple benchmarks [94, 93].
IV-D Generalized Video Alignment
The final step in meta-learning-based FSAR is inference. This stage typically involves comparing query samples to class prototypes and assigning the most likely class based on similarity measures. However, due to the temporal nature of video features, alignment is complex. Several studies perform fine-grained alignment at the frame-level rather than treating the video as a whole, emphasizing long-term temporal ordering information to ensure consistency and accuracy of matching. In addition to temporal information, appearance features, global features, and semantic features contained in the video are also crucial. Therefore, several methods propose multi-level alignment to simultaneously focus on different dimensions of video features, thereby achieving more reliable matching.
IV-D1 Temporal Alignment
@@
Several methods propose to highlight advancements in temporal alignment for video similarity measurement, balancing strengths in accuracy and adaptability with challenges in computational complexity and parameter tuning. For instance, to simplify the sequence matching procedure, ITANet [70] learns an implicit temporal alignment that adopts a self-attention mechanism to augment frame features with their temporal context, reducing temporal variation in video instances. OTAM [69] pioneers the utilization of temporal ordering information by integrating segment distances solely along the ordered temporal alignment path which is illustrated in Figure 7(c). This method extends the Dynamic Time Warping (DTW) algorithm [142] to enforce the distance for prediction to preserve temporal ordering. Moreover, its differentiable nature enhances flexibility and effectiveness in utilizing long-term temporal information. Li et al. [143] introduce event boundaries into FSAR to serve as a guiding prior for temporal alignment. The authors uniformly sample frames from each action subpart based on these boundaries, rather than considering the entire video. This frame sampling strategy is rooted in temporal boundaries and devised to alleviate intra-class variance by ensuring uniform representation across action segments. Furthermore, a boundary selection module is introduced to pinpoint the initiation and conclusion of actions, facilitating precise alignment of videos with the respective durations. TRX [84] performs segment-to-segment alignment (Figure 7(d)), aligning temporally-ordered sub-sequences rather than individual frames to effectively match actions performed at varying speeds and in different parts of videos. Wang et al. [115] develop a motion-augmented long-short contrastive learning method to jointly model the global contextual information and motion dynamics. To explicitly integrate the global context into the local matching process, a long-short contrastive objective is applied to enforce frame features to predict the global context of the videos that belong to the same class. The authors take frame differences as motion information and design a motion autodecoder to explicitly extract motion features between frame representations by reconstructing pixel motions. Li et al. [144] adopt the Earth Mover’s Distance in the transportation problem to measure the similarity between video features in terms of sub-action representations. It computes the optimal matching flows between sub-actions as a distance metric, which is favorable for comparing fine-grained patterns.
@@
IV-D2 Multi-level Alignment
@@
IV-D3 Discussion
Achieving effective video alignment in few-shot action recognition is challenging due to the presence of fine-grained motion changes, the temporal nature of videos, and irrelevant sub-movements that can cause misalignment. To address these issues, existing methods have made significant contributions, as demonstrated by their performance on common benchmarks (Table IV). Most current approaches emphasize the acquisition and utilization of temporal information during the alignment, which is crucial for the effectiveness of the matching methods. Additionally, the robustness of matching can be further enhanced by integrating multi-level alignment techniques.
TABLE IV: Results of different Generalized Video Alignment methods.
@@ V Advanced Topics
V-A Skeleton-based Learning
@@
In Section IV, we summarize the contributions of existing few-shot action recognition works which are primarily based on commonly available RGB modalities. However, beyond RGB, human actions can be captured using diverse data modalities that encode unique information, each offering distinct advantages tailored to specific application scenarios. Skeleton data is a highly effective modality for identifying human behavior, finding extensive applications in traditional behavior recognition tasks. Several research efforts aim to address the problem of skeleton identification in few-shot scenarios [155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167]. Sabater et al. [155] introduce skeleton-based representations into FSAR, computing motion representations robust to variable kinematic conditions. This approach involves evaluating query and target motion representations. Specifically, the query action is processed and compared against a target motion sequence to recognize when the anchor action has been performed within the target motion sequence. Ma et al. [157] introduce a spatial matching strategy comprising spatial disentanglement and spatial activation. The rationale behind spatial disentanglement stems from the observation that augmenting spatial information for leaf nodes (e.g., the ’hand’ joint) enhances representation diversity for skeleton matching. To realize spatial disentanglement, the authors advocate for the representation of skeletons in a full-rank space with a rank maximization constraint. Furthermore, they propose an attention-based spatial activation mechanism to incorporate disentanglement, whereby the disentangled joints are adaptively adjusted based on matching pairs. PAINet [161] integrates two complementary branches aimed at enhancing inter-skeleton and intra-skeleton correlations by leveraging topology and physical information to refine the modeling of interactive parts and joint pairs within both branches. Moreover, a spatial cross-attention module is utilized to establish joint associations across sequences, while introducing a directional Average Symmetric Surface Metric to identify the closest temporal similarity. Concurrently, a spatial self-attention mechanism aggregates spatial context within sequences, and a temporal cross-attention network is applied to correct temporal misalignment and calculate similarity. Li et al. [158] propose a self and mutual adaptive matching module designed to transform these feature maps into more discriminative feature vectors. Initially, it exploits both the temporal information inherent in each individual skeleton joint and the spatial relationships among them to extract features. Subsequently, the SMAM module dynamically assesses the similarity between labeled and query samples, facilitating feature matching within the query set to differentiate between similar skeletons across various action categories. Wei et al. [164] propose leveraging a contrastive diffusion graph convolutional network (CD-GCN) for few-shot action recognition. Their approach relies on attentional diffusion and contrastive loss to generate robust representations, facilitating the diffusion of diverse graph semantic features. Skeleton data inherently provides straightforward yet informative representations of body structure and pose information. These skeleton-based methodologies demonstrate robustness against variations in action-independent contexts and have garnered increasing attention.
V-B Multimodal Learning
@@
Because videos have the natural multimodal characteristic, several methods [98, 83, 112, 168, 169, 170, 62, 171, 114, 172, 173] propose to leverage multimodal information (e.g. optical flow, depth, audio and compressed domain data) to enhance the few-shot action recognition procedure. For example, AMFAR [172] proposes to actively find the reliable modality, either RGB or optical flow, for each sample based on task-dependent context information to improve the few-shot reasoning procedure. Additionally, optical flow, depth, and appearance features are leveraged in Lite-MKD [173] to train a teacher model to achieve a comprehensive fusion of human movements. A student model is utilized to learn to recognize actions from the single RGB modality at a lower computational cost under the guidance of the teacher. Mercea et al. [174] introduce a unified audio-visual few-shot video classification benchmark. They propose a text-to-feature diffusion framework that fuses temporal and audio-visual features through cross-modal attention and generates multi-modal features for novel classes. This framework incorporates a hybrid attention mechanism and a diffusion model for multi-modal feature generation, thereby expanding the training dataset with additional novel-class samples.
@@ V-C Unsupervised Learning
V-D Cross-Domain
Current few-shot action recognition methods assume abundant labeled videos of base classes and few-shot videos of novel classes originating from the same domain. However, this assumption is impractical as novel categories often emerge from different data domains with diverse spatial and temporal characteristics in real-world applications. To tackle this issue, Gao et al. [175] propose investigating cross-domain few-shot action recognition. They employ an attentive adversarial network to align spatiotemporal dynamics, addressing domain discrepancies and enhancing the discrimination of learned domain-invariant features through a pairwise margin discrimination loss. CDFSL-V [176] utilizes self-supervised learning and curriculum learning to balance information from both the source and target domains, thereby learning generic features for cross-domain few-shot action recognition. Li et al. [177] introduce a method that identifies and leverages transferable knowledge from a model pretrained on a large image dataset. During adaptation, this approach preserves transferable temporal dynamics while updating the image encoder. Additionally, SSAlign [178] and RelaMix [179] explore few-shot domain adaptation strategies that diverge from the conventional definition of FSAR proposed in this paper but inspire further investigation into adapting models across domains. Cross-domain few-shot action recognition addresses the challenge of recognizing actions in novel domains with limited labeled data, which is crucial for applications requiring the adaptation of existing models to new environments or scenarios.
V-E Incremental Learning
@@ V-F Federated Learning
With the widespread use of action recognition technology, ensuring data privacy and addressing ethical concerns have become increasingly critical. This involves making the decision-making process traceable, thereby bolstering user trust in the technology. In few-shot action recognition, this necessitates the development of privacy-preserving methods such as differential privacy and federated learning to enable effective action recognition while safeguarding personal data. Federated learning is a promising solution for many computer vision applications due to its effectiveness in handling data privacy and communication overhead. For example, Tu et al. [185] propose investigating the federated few-shot learning problem, which involves collaboratively learning a classification model from multiple FL clients to recognize unseen actions with only a few labeled video samples.
VI Future work
VI-A More Realistic Dataset

### v2

@@ 4.2.3 Discussion
TSAM [67] ArXiv’2024 CLIP-ViT-B 65.8 74.6 96.2 97.1 98.3 99.3 60.5 66.7 84.5 88.9
MA-FSAR [176] ArXiv’2024 CLIP-ViT-B 63.3 72.3 95.7 96.0 97.2 99.2 59.1 64.5 95.7 96.0
Task-Adapter [12] ACM MM’2024 CLIP-ViT-B 71.3 74.2 95.0 96.8 98.0 99.4 60.2 70.2 83.6 88.8
Manta [52] ACM MM’2024 ImageNet-VM 66.1 89.1 84.4 96.2 96.9 99.4 - - 89.1 96.6
@@
4.3 Category Prototype Learning
@@ 4.3.1 Aggregation-based
Text-Guided. Furthermore, several methods are inspired by the strong generalization ability of textual information to modulate the action prototypes with the guidance of semantic priors to resolve the challenge of inaccurate prototype estimation caused by data scarcity [99, 175], as illustrated in Figure 6(d). For instance, Ni et al. [99] extract semantic features from label texts with a frozen CLIP [108] text encoder and further enhanced by a semantic-enhanced module. The prototypes are enhanced by semantic features with flexible techniques such as weighted average and multi-head attention. Similarly, Wang et al. [159] adaptively refine visual prototypes with transferable textual concepts from CLIP, which is achieved by implementing a temporal Transformer to fuse the textual and visual features in the support set. These approaches demonstrate how integrating multimodal semantic information can significantly improve the fidelity and effectiveness of prototype representations.
4.3.2 Learning-based
@@
4.3.3 Compound Methods
Due to the intertwining of temporal and spatial information in videos, alongside the potential occurrence of actions at diverse speeds, a single approach to prototype calculation may overlook essential video details. Consequently, certain methodologies advocate for computing compound prototypes to fortify the resilience of prototype representations. For instance, Huang et al. [53] summarize each video into compound prototypes consisting of a group of global prototypes and a group of focused prototypes, and then compare video similarity based on the compound prototypes. Each global prototype summarizes a specific aspect of the entire video, while focused prototypes target certain timestamps, with bipartite matching applied to handle temporal variations in actions. The authors further introduce a method to learn prototypes combined with three types of prototypes [54]. The first type describes specific aspects of the action regardless of its timestamp. The second type focuses on specific timestamps and uses bipartite matching to handle temporal variations. The third type is generated from the timestamp-centered prototypes, regularizing temporal consistency and providing an auxiliary summarization of the entire video. Jiang et al. [58] introduce a dual-prototype network that integrates class-specific and query-specific attentive learning. The class-specific attentive learning approach enhances prototype representativeness by calculating within-class similarity for each support sample, thereby reducing the influence of noise and outliers. In contrast, the query-specific attention mechanism adaptively reweights support samples based on their similarities to each query sample, which increases inter-class discrimination by enlarging differences among support samples. Guo et al. [42] propose a Dual Motion-Guided Attention Learning method to efficiently and effectively identify and correlate motion-related region features from the video level to the task level, allowing the model to construct class prototypes that fully incorporate spatio-temporal relationships. Wu et al. [169] introduce a multi-level matching algorithm that involves comparing visual information and textual information, as well as comparing the visual information conveyed by the support and query samples respectively. In summary, these methodologies highlight the effectiveness of compound prototype representations in bolstering prototype robustness significantly and enhancing the ability to generalize to novel categories.
4.3.4 Discussion
Table 3: Results of different category prototype learning methods.
@@ 4.3.4 Discussion
DMGAL-FT [42] ArXiv’2024 Resnet-50 55.5 71.1 72.5 88.0 87.1 97.4 41.2 61.8 60.2 79.3
EMP-Net [169] ECCV’2024 CLIP-ViT-B 63.1 73.0 89.1 93.5 94.3 98.2 57.1 65.7 76.8 85.8
The objective of category prototype learning in few-shot action recognition is to learn representative and generalized action representations for novel categories from a few samples, thereby achieving stable recognition performance in each meta-task. The conventional average-based method may be affected by sample capacity and extreme samples, making mean-of-class prototypes inadequate in representing the distribution of samples. As shown in Table 3, the prototype learning methods designed for FSAR have achieved significant performance. More recent approaches that enhance class prototype representations with semantic prior knowledge of language image pretrained models have demonstrated significant performance benefits across multiple benchmarks [99, 159].
4.4 Generalized Video Alignment
The final step in meta-learning-based FSAR is inference. This stage typically involves comparing query samples to class prototypes and assigning the most likely class based on similarity measures. However, due to the temporal nature of video features, alignment is complex. Several studies perform fine-grained alignment at the frame level rather than treating the video as a whole, emphasizing long-term temporal ordering information to ensure consistency and accuracy of matching. In addition to temporal information, appearance features, global features, and semantic features contained in the video are also crucial. Therefore, several methods propose multi-level alignment to simultaneously focus on different dimensions of video features, thereby achieving more reliable matching.
4.4.1 Temporal Alignment
@@
Several methods propose to highlight advancements in temporal alignment for video similarity measurement, balancing strengths in accuracy and adaptability with challenges in computational complexity and parameter tuning. For instance, to simplify the sequence matching procedure, ITANet [193] learns an implicit temporal alignment that adopts a self-attention mechanism to augment frame features with their temporal context, reducing temporal variation in video instances. OTAM [13] pioneers the utilization of temporal ordering information by integrating segment distances solely along the ordered temporal alignment path which is illustrated in Figure 7(c). This method extends the Dynamic Time Warping (DTW) algorithm [97] to enforce the distance for prediction to preserve temporal ordering. Moreover, its differentiable nature enhances flexibility and effectiveness in utilizing long-term temporal information. Li et al. [70] introduce event boundaries into FSAR to serve as a guiding prior for temporal alignment. The authors uniformly sample frames from each action subpart based on these boundaries, rather than considering the entire video. This frame sampling strategy is rooted in temporal boundaries and devised to alleviate intra-class variance by ensuring uniform representation across action segments. Furthermore, a boundary selection module is introduced to pinpoint the initiation and conclusion of actions, facilitating precise alignment of videos with the respective durations. TRX [104] performs segment-to-segment alignment (Figure 7(d)), aligning temporally-ordered sub-sequences rather than individual frames to effectively match actions performed at varying speeds and in different parts of videos. Wang et al. [160] develop a motion-augmented long-short contrastive learning method to jointly model the global contextual information and motion dynamics. To explicitly integrate the global context into the local matching process, a long-short contrastive objective is applied to enforce frame features to predict the global context of the videos that belong to the same class. The authors take frame differences as motion information and design a motion autoencoder to explicitly extract motion features between frame representations by reconstructing pixel motions. Li et al. [68] adopt the Earth Mover’s Distance in the transportation problem to measure the similarity between video features in terms of sub-action representations. It computes the optimal matching flows between sub-actions as a distance metric, which is favorable for comparing fine-grained patterns.
@@
4.4.2 Multi-level Alignment
@@
4.4.3 Discussion
Achieving effective video alignment in few-shot action recognition is challenging due to the presence of fine-grained motion changes, the temporal nature of videos, and irrelevant sub-movements that can cause misalignment. To address these issues, existing methods have made significant contributions, as demonstrated by their performance on common benchmarks (Table 4). Most current approaches emphasize the acquisition and utilization of temporal information during the alignment, which is crucial for the effectiveness of the matching methods. Additionally, the robustness of matching can be further enhanced by integrating multi-level alignment techniques.
Table 4: Results of different generalized video alignment methods.
@@ 5 Advanced Topics
5.1 Skeleton-based Learning
@@
In Section 4, we summarize the contributions of existing few-shot action recognition works which are primarily based on commonly available RGB modalities. However, beyond RGB, human actions can be captured using diverse data modalities that encode unique information, each offering distinct advantages tailored to specific application scenarios. Skeleton data is a highly effective modality for identifying human behavior, finding extensive applications in traditional behavior recognition tasks. Several research efforts aim to address the problem of skeleton identification in few-shot scenarios [111, 151, 89, 74, 139, 92, 83, 168, 180, 167, 86, 32, 185]. Sabater et al. [111] introduce skeleton-based representations into FSAR, computing motion representations robust to variable kinematic conditions. This approach involves evaluating query and target motion representations. Specifically, the query action is processed and compared against a target motion sequence to recognize when the anchor action has been performed within the target motion sequence. Ma et al. [89] introduce a spatial matching strategy comprising spatial disentanglement and spatial activation. The rationale behind spatial disentanglement stems from the observation that augmenting spatial information for leaf nodes (e.g., the ’hand’ joint) enhances representation diversity for skeleton matching. To realize spatial disentanglement, the authors advocate for the representation of skeletons in a full-rank space with a rank maximization constraint. Furthermore, they propose an attention-based spatial activation mechanism to incorporate disentanglement, whereby the disentangled joints are adaptively adjusted based on matching pairs. PAINet [83] integrates two complementary branches aimed at enhancing inter-skeleton and intra-skeleton correlations by leveraging topology and physical information to refine the modeling of interactive parts and joint pairs within both branches. Moreover, a spatial cross-attention module is utilized to establish joint associations across sequences, while introducing a directional Average Symmetric Surface Metric to identify the closest temporal similarity. Concurrently, a spatial self-attention mechanism aggregates spatial context within sequences, and a temporal cross-attention network is applied to correct temporal misalignment and calculate similarity. Li et al. [74] propose a self and mutual adaptive matching module designed to transform these feature maps into more discriminative feature vectors. Initially, it exploits both the temporal information inherent in each individual skeleton joint and the spatial relationships among them to extract features. Subsequently, the SMAM module dynamically assesses the similarity between labeled and query samples, facilitating feature matching within the query set to differentiate between similar skeletons across various action categories. Wei et al. [167] propose leveraging a contrastive diffusion graph convolutional network (CD-GCN) for few-shot action recognition. Their approach relies on attentional diffusion and contrastive loss to generate robust representations, facilitating the diffusion of diverse graph semantic features. Skeleton data inherently provides straightforward yet informative representations of body structure and pose information. These skeleton-based methodologies demonstrate robustness against variations in action-independent contexts and have garnered increasing attention.
5.2 Multimodal Learning
@@
Because videos have the natural multimodal characteristic, several methods [179, 7, 31, 47, 45, 64, 164, 90, 87, 166, 78] propose to leverage multimodal information (e.g. optical flow, depth, audio and compressed domain data) to enhance the few-shot action recognition procedure. For example, AMFAR [166] proposes to actively find the reliable modality, either RGB or optical flow, for each sample based on task-dependent context information to improve the few-shot reasoning procedure. Additionally, optical flow, depth, and appearance features are leveraged in Lite-MKD [78] to train a teacher model to achieve a comprehensive fusion of human movements. A student model is utilized to learn to recognize actions from the single RGB modality at a lower computational cost under the guidance of the teacher. Mercea et al. [93] introduce a unified audio-visual few-shot video classification benchmark. They propose a text-to-feature diffusion framework that fuses temporal and audio-visual features through cross-modal attention and generates multi-modal features for novel classes. This framework incorporates a hybrid attention mechanism and a diffusion model for multi-modal feature generation, thereby expanding the training dataset with additional novel-class samples. With the emergence of visual language pre-training models, endeavors have been made to leverage the textual information of action labels to enhance the accuracy of few-shot action recognition FSAR [159, 99, 12, 169, 176].
@@ 5.3 Unsupervised Learning
5.4 Cross-Domain
Current few-shot action recognition methods assume abundant labeled videos of base classes and few-shot videos of novel classes originating from the same domain. However, this assumption is impractical as novel categories often emerge from different data domains with diverse spatial and temporal characteristics in real-world applications [34, 113, 91, 41, 165]. To tackle this issue, Gao et al. [34] propose investigating cross-domain few-shot action recognition. They employ an attentive adversarial network to align spatiotemporal dynamics, addressing domain discrepancies and enhancing the discrimination of learned domain-invariant features through a pairwise margin discrimination loss. CDFSL-V [113] utilizes self-supervised learning and curriculum learning to balance information from both the source and target domains, thereby learning generic features for cross-domain few-shot action recognition. Li et al. [73] introduce a method that identifies and leverages transferable knowledge from a model pretrained on a large image dataset. During adaptation, this approach preserves transferable temporal dynamics while updating the image encoder. Additionally, SSAlign [181] and RelaMix [102] explore few-shot domain adaptation strategies that diverge from the conventional definition of FSAR proposed in this paper but inspire further investigation into adapting models across domains. Cross-domain few-shot action recognition addresses the challenge of recognizing actions in novel domains with limited labeled data, which is crucial for applications requiring the adaptation of existing models to new environments or scenarios.
5.5 Incremental Learning
@@ 5.6 Federated Learning
With the widespread use of action recognition technology, ensuring data privacy and addressing ethical concerns have become increasingly critical. This involves making the decision-making process traceable, thereby bolstering user trust in the technology. In few-shot action recognition, this necessitates the development of privacy-preserving methods such as differential privacy and federated learning to enable effective action recognition while safeguarding personal data. Federated Learning (FL) is a promising solution for many computer vision applications due to its effectiveness in handling data privacy and communication overhead. For example, Tu et al. [140] propose investigating the federated few-shot learning problem, which involves collaboratively learning a classification model from multiple FL clients to recognize unseen actions with only a few labeled video samples.
6 Future work
6.1 More Realistic Dataset